Suppose {an} ∞ n=1 be a sequence of positive real numbers and 0 < x < 1. If an+1 < x · an for every n ∈ N, prove that limn→∞ an = 0.


1
Expert's answer
2021-02-24T07:31:51-0500

First of all,

By applying the inequality multiple times


we find that "a_n < x^na_0, \forall n \in \mathbb{N}" (as "a_n < x a_{n-1}<x^2a_{n-2}<...<x^n a_0").


Now as "a_n \geq 0, \forall n\in \mathbb{N}", and for any "a_0 \in \mathbb{R}^+", "a_0x^n \underset{n\to\infty}{\to}0", as "0<x<1",


We see that "0 \leq a_n \leq a_0 x^n", the right sequence converges to zero and thus we can conclude


that "a_n \underset{n\to\infty}{\to} 0" as it is bounded by two sequences both converging to zero.